OpenStudy (anonymous):

What polynomial identity should be used to prove that 133 = 125 + 8? Difference of Cubes Difference of Squares Square of Binomial Sum of Cubes

OpenStudy (mathmate):

In general, \(a^3+b^3=(a+b)(a^2-ab+b^2)\) \(a^3-b^3=(a-b)(a^2+ab+b^2)\) Specifically for this case, \(125=5^3; 8=2^3; so 125+8=5^3+2^3=...\)

OpenStudy (anonymous):

im thinking it's difference of cubes

OpenStudy (anonymous):

you still there

OpenStudy (mathmate):

hint: sum is the result of addition difference is the result of subtraction.